# An effective control of large systems of active particles: An application to evacuation problem

### v1

@@ Abstract
[sk] organization=Applied AI Center, Skolkovo Institute of Science and Technology, addressline=Bolshoy Boulevard, 30, bld.1, city=Moscow, postcode=121205, country=Russia
@@ 2.1 General statement of the evacuation problem
The main limitation of such models is the computational complexity, associated with the complete implementation of the SFM for a large system of agents. Therefore the lattice-automata approach, was widely used, with the set of discrete actions and states. Here the pedestrians are individual ”particles”, with an appropriate speed and size; their movement is governed by spatial transition probabilities. This allows encoding of simple rules of behavior in numerical simulations, which is realized in a number of available software Legion [40], EXITUS [30], Simulex [41], EXIT89 [42], PEDFLOW [43], etc.)
@@ 2.2 Generalized Vicsek Model
where $\vec{r_{i}}(t)$ and $\vec{r_{i}}(t+\Delta t)$ are respectively the coordinate of $i$-th particle (in the ensemble of $N$ particles) at time $t$ and $t+\Delta t$. Each particle has a velocity $\vec{v}_{i}(t)=v\vec{e}_{i}(t)$ at time $t$, which absolute value is $v$ and the direction is determined by the unit vector $\vec{e_{i}}\texttt{(t) = }\left(\cos\theta_{i}\texttt{(t)},\sin\theta_{i}\texttt{(t)}\right)^{T}$, with the azimuthal angle $\theta_{i}$. The radius of neighbour influence is $r$, and $\zeta_{i}(t)$ is a random value, corresponding to the noise. In the original Vicsek model it is a random number with a uniform probability from the interval $\eta/2,\eta/2]$, where $\eta$ is the noise magnitude. The noise covariance reads, $\langle\zeta_{i}\zeta_{j}\rangle=\sigma^{2}\delta_{ij}$, with the dispersion $\sigma$.
@@ 2.3 Evacuation Simulation
is the unit vector for $i$-th particle to the exit destination. At the end, having reached the evacuation destination – the small circle with an exit point at the center, Figure 2, an individual is considered saved. For clear simulations and training, the environment is implemented [^2] Source is available on GitHub: https://github.com/cinemere/evacuation in style of OpenAI Gym environment ( https://www.gymlibrary.ml/)[56].
@@ 3.2 Training details
In the present work, we compare three methods: feed-forward, transformer, and a novel approach based on pseudo-gravitational encoding. The primary distinction among these methods lies in their neural network architectures. It is important to note that the pseudo-gravitational encoding is not a standalone neural architecture, but rather a state-representation scheme that acts as the first layer of the network. The encoded forces are then passed to the same feed-forward network architecture as in the baseline model, with the only difference being the input representation. To train the action policy, we utilized the Robust Policy Optimization (RPO) [63] implementation from the CleanRL [64] package. More details on the RL algorithm are presented in A.
@@
    $T_{j}=(\vec{r}_{j}-\vec{r}_{l},\,i_{j})$ for $j=1..N$, where $i_{j}$ is one-hot encoding whether pedestrian $j$ is in the Vicsek zone, the catch zone, the exit zone, or already evacuated.
@@ 4 Experimental Results
Figure 4: The curves, representing the training process: the average episode reward, as a function of the simulation time step. The results are averaged over 10 training runs with different environment random seeds; the shaded areas depicts the the standard error. The green curve corresponds to RL with the “pseudo-gravitational” encoding, orange - RL with Transformer architecture and blue – RL with Feed Forward architecture. The exponent of the “pseudo-gravitational” potential $\alpha=1$.
@@ Appendix B Average Angle/Direction
then the average angle is defined as follows [^6] The relations below correspond to the standard function atan2($p$).
@@
and corresponds to the standard formulation of the Vicsek model [5]. We use the angle notation here for compactness only. It is important to stress that a simple arithmetic mean of angles is not well-defined, since its result depends on the chosen branch of the angle interval and may yield contradictory outcomes. By contrast, averaging in the vector form and then computing the argument ensures a unique definition of the mean direction. In degenerate symmetric cases (when the average vector has zero magnitude), no prevailing orientation exists and the new direction is undefined what is also reflected in the equation.
@@ Appendix D Parameters of the Model and “Pseudo-Gravitational” Potential
$L$, units Height & width of simulation area 11

### v2

@@ Abstract
[sk] organization=Artificial Intelligence Center, Skolkovo Institute of Science and Technology, addressline=Bolshoy Boulevard, 30, bld.1, city=Moscow, postcode=121205, country=Russia
@@ 2.1 General statement of the evacuation problem
The main limitation of such models is the computational complexity, associated with the complete implementation of the SFM for a large system of agents. Therefore the lattice-automata approach, was widely used, with the set of discrete actions and states. Here the pedestrians are individual ”particles”, with an appropriate speed and size; their movement is governed by spatial transition probabilities. This allows encoding of simple rules of behavior in numerical simulations, which is realized in a number of available software ( Legion [40], EXITUS [30], Simulex [41], EXIT89 [42], PEDFLOW [43], etc.)
@@ 2.2 Generalized Vicsek model
where $\vec{r_{i}}(t)$ and $\vec{r_{i}}(t+\Delta t)$ are respectively the coordinate of $i$-th particle (in the ensemble of $N$ particles) at time $t$ and $t+\Delta t$. Each particle has a velocity $\vec{v}_{i}(t)=v\vec{e}_{i}(t)$ at time $t$, which absolute value is $v$ and the direction is determined by the unit vector $\vec{e_{i}}\texttt{(t) = }\left(\cos\theta_{i}\texttt{(t)},\sin\theta_{i}\texttt{(t)}\right)^{T}$, with the azimuthal angle $\theta_{i}$. The radius of neighbour influence is $r$, and $\zeta_{i}(t)$ is a random value, corresponding to the noise. In the original Vicsek model it is a random number with a uniform probability from the interval $\eta/2,\eta/2]$, where $\eta$ is the noise magnitude.
@@ 2.3 Evacuation simulation
is the unit vector for $i$-th particle to the exit destination. At the end, having reached the evacuation destination – the small circle with an exit point at the center, Figure 2, an individual is considered saved. For clear simulations and training, the environment has been implemented with the interface of OpenAI Gym environment [^2] Source is available on GitHub: https://github.com/cinemere/evacuation with the interface of OpenAI Gym environment (https://gymnasium.farama.org)[56]..
@@ 3.2 Training details
In the present work, we compare three methods: feed-forward, transformer, and a novel approach based on pseudo-gravitational encoding. The primary distinction among these methods lies in their neural network architectures. It is important to note that the pseudo-gravitational encoding is not a standalone neural architecture, but rather a state-representation scheme that acts as the first layer of the network. The encoded forces are then passed to the same feed-forward network architecture as in the baseline model, with the only difference being the input representation. To train the action policy, we utilized the Robust Policy Optimization (RPO) [63] implementation from the CleanRL [64] package. The complete set of RPO agent hyperparameters is summarized in Table 4, while additional details on the RL algorithm are presented in A.
@@
    $T_{j}=(\vec{r}_{j}-\vec{r}_{l},\,i_{j})$ for j=1Nj=1...N, where $i_{j}$ is one-hot encoding whether pedestrian $j$ is in the Vicsek zone, the catch zone, the exit zone, or already evacuated.
@@ 4 Experimental results
Figure 4: The curves, representing the training process: the average episode reward, as a function of the simulation time step. The results are averaged over 10 training runs with different environment random seeds; the shaded areas depicts the standard error. The green curve corresponds to RL with the “pseudo-gravitational” encoding, orange - RL with Transformer architecture and blue – RL with Feed Forward architecture. The exponent of the “pseudo-gravitational” potential $\alpha=1$.
@@ Appendix B Average angle/direction
then the average angle is defined as follows [^6] The relations below correspond to the standard function atan2($y_{p},x_{p}$).
@@
and corresponds to the standard formulation of the Vicsek model [5]. We use the angle notation here for compactness only. It is important to stress that a simple arithmetic mean of angles is not well-defined, since its result depends on the chosen branch of the angle interval and may yield contradictory outcomes. By contrast, averaging in the vector form and then computing the argument ensures a unique definition of the mean direction. In degenerate symmetric cases (when the average vector has zero magnitude), no prevailing orientation exists and the new direction is undefined which is also reflected in the equation.
@@ Appendix D Parameters of the model and “Pseudo-Gravitational” potential
$L$, units Square simulation domain: [L,L][-L,L] in both $x$ and $y$ directions 11